Two cars leave town at the same time going in the opposite direction. One travels 35 mph and the other travels 49 mph. In how many hours will they be 126 miles apart?
Answer by rwm(914)

r*t=d
35t+49t=126
84t=126
42t=63
6t=9
t=9/6
t=1.5 hours
35*1.5+49*1.5=126